A particle executes a simple unrestricted random walk on the line: a step to the right of length 1 occurs with probability 0.66 and a step to the left of length 1 occurs with probability 0.34. Initially the particle is at the origin.

Calculate the probability that the particle will ever return to the origin.
Answer #1

P(right) = p = 0.66

P(left) = q = 0.34

The particle will return to origin if the number of right and left are equal.

So, P(return to origin) = (1 right, 1 left)+(2 right, 2 left)+...

= pq+p2q2+p3q3+....

= pq/(1-pq) (sum to infinite terms in G.P)

= (0.66*0.34)/(1-(0.66*0.34)) = 0.2244/(1-0.2244) = 0.2893